A house is flown with a 100watts bulb, a 1,000watts, a 40watts radio and a 200watts bulb. If the house is supplied with 200volts. Find the current passing through the main fuse when all appliances are ON.


1
Expert's answer
2022-02-24T11:30:49-0500

Answer

Total consumed power

P=100+1000+40+200=1340W

Voltage supplied

V=200V

Current passing

"I=\frac{P}{V}\\=\frac{1340}{200}\\=11.7A"